### v1

# Pulse-induced memory-like effect in cyclotron motion?

Abstract
We study how a charged particle moving in a uniform magnetic field along its standard circular path (cyclotron motion) reacts to a short-duration, homogeneous, uniform electric field pulse injected in the plane perpendicular to the magnetic field. A ‘permanent’ change in the radius of the initial circle and a shift of its centre is noted at later times, after the pulse is switched off. The magnitude of the velocity undergoes a change too, akin to a ‘velocity kick’. In summary, our results suggest a pulse-induced ‘electromagnetic memory-like effect’, which is not quite a ‘wave memory’, but, nevertheless, has similar features within a simple, non-relativistic context.
I Introduction
Gravitational wave memory–a ‘permanent change’ caused by a gravitational wave pulse–has been a topic of active research interest over the last few years. Though known to us since the papers of Zel’dovich, Polnarev zelpol and Braginsky, Grishchuk bragri and subsequent seminal work by many authors (eg. christo ), the recent upsurge in research is largely motivated by prospects of observations in the context of gravitational wave physics memoryexpt0 ; memoryexpt1 ; memoryexpt2 . Different aspects of gravitational wave memory have been analysed in a large number of papers biegarf ; tolwal ; tolwal1 ; madwin1 ; madwin2 ; stro1 ; stro2 ; zhang1 ; zhang2 ; zhang3 ; zhang4 ; zhang5 ; ic1 ; cvet ; ic2 ; ic3 ; ic4 ; pah1 ; pah2 ; jpuzan ; eanna1 ; eanna2 ; eanna3 ; olough ; diva ; shore ; mota . Reviews on gravitational wave memory are available in favata1 , cqgfocus1 .
In analogy with gravitational wave memory, electromagnetic memory is conventionally defined also through a ‘permanent change’– the so-called ‘velocity kick’. It was first briefly discussed in a paper by Grishchuk and Polnarev grishpol and has recently been extensively analysed by Bieri and Garfinkle biegarf1 . It is defined (for a specific case where the force is $q{\bf E}$ and unit mass) using the following simple equation:
[…]
where ${\bf E}$ is the electric field, usually taken as a radiation field due to a source far away. In the above formula, if one considers an electric field which is non-zero and constant only over a small interval (0,T) in time, then the R. H. S. simply becomes ${q\bf E_{0}}T$ and we have the velocity at negative infinity (past) having a different value compared to its value at positive infinity (future). This also implies a jump or a kick in the velocity, imparted to the charged test particle by the electric pulse. Conventionally though, one considers the electric field due to radiation–say, electric dipole radiation (far away from the source) – and its effect on a test charge. In the non-relativistic case, one can evaluate the time rate of change of the dipole moment of the source in the past and future and relate it to the charges and velocities of the source. This would give an expression for the velocity kick on the test charge in terms of the behaviour of the source in the past and future biegarf2 . The analysis is somewhat more intricate when one moves away from the slow motion approximation for the source entities. It is available in biegarf2 . The essence however is to see if the velocity of the test charge changes on account of the field or the pulse (i.e. we find a nonzero value of the L. H. S. in Eqn. (1)), and whether there is a ‘permanent change’ which appears after the pulse has departed, and persists later too. It is this viewpoint, which does not necessarily attribute memory to a passing wave field, is what we focus on in this article. Diverse aspects of electromagnetic memory as well as ‘color’ (Yang-Mills) memory have been discussed in numerous papers in the recent past winicour ; pasterski ; pate ; satish ; andrea ; garfinkle ; biegarf2 ; seraj .
In the example which we discuss here, we will consider a particle of charge $q$ and mass $m$ placed in a uniform magnetic field and encountering a spatially homogeneous electric pulse. Prior to the arrival of the pulse, the particle travels in a circular path. After the pulse has departed, what happens to this trajectory? How is it influenced by the pulse? Is that influence a permanent change? These questions when answered (as we do below) will shed light on an effect quite similar to ‘electromagnetic memory’ , albeit in a rather simple, purely non-relativistic context. We emphasize here, that we are not considering a ‘wave’ memory–rather we are defining ‘memory’ arising out of a short duration electric pulse in a background uniform magnetic field. From our perspective, the memory aspect is largely encoded in a ‘persistent and permanent change’ caused by the pulse and we are interested in quantifying this change in our work.
II Charged particle in a uniform magnetic field and an electric pulse
Let us now state our problem more quantitatively and obtain a solution. We shall assume the magnetic field to be uniform and along the $z$ direction, i.e. ${\bf B}=B{\bf\hat{k}}$. The electric field is in the $xy$ plane and we assume it to be:
${\bf E}=E(t)\left(\cos\,\alpha{\bf\hat{i}}+\sin\,\alpha{\bf\hat{j}}\right)$ (2)
where $E$ is chosen as the following function (see Figure 1):
$E(t)=0,\hskip 28.90755pt-\infty\leq t\leq 0$ (3)
$=E,\hskip 36.135pt0\leq t\leq T$
$=0\hskip 65.04256ptt\geq T$
with the ‘E’ introduced above, a constant.
Figure 1: The square pulse $E$ as function of $t$.
This is the well-known square pulse, as shown in Figure 1.
In the standard cyclotron, the electric field is chosen to be an alternating harmonic field applied between the dees and oscillating at a frequency in the radio-frequency range. This leads to a spiral path and the charged particle gets accelerated in the process, thereby achieving the aim of an accelerator chautard . In contrast here we apply a short duration electric field with the aim of injecting a visible difference between trajectories before and after the duration of the pulse.
The equations of motion for our system in the interval $(0,T)$ (where the electric pulse acts) are easy to write down ($m\frac{d{\bf v}}{dt}=q{\bf E}+q{\bf v}\times{\bf B}$) and are given, in component form, as:
$m\frac{dv_{x}}{dt}=qE\cos\alpha+qv_{y}B$ (4)
$m\frac{dv_{y}}{dt}=qE\sin\alpha-qv_{x}B$ (5)
$m\frac{dv_{z}}{dt}=0$ (6)
where $q$ and $m$ are the charge and mass of the particle, respectively, and ${\bf v}=\left(v_{x},v_{y},v_{z}\right)$ is its velocity vector. In the other two regions, i.e. for $t\leq 0$ and $t\geq T$, the same equations hold with $E=0$. The $v_{z}$ equation is trivially solved and we shall assume $v_{z}=0$ by setting the integration constant to zero, thereby restricting the motion of the charge to the $xy$ plane.
The well-known way to solve the equations for $v_{x}$ and $v_{y}$ is to differentiate one of them first and use the other equation. This will give a second order uncoupled equation for $v_{x}$ (or $v_{y}$). One can then use the solution of this second order equation to find the other velocity component using one of the first order equations.
For the problem at hand, we will need to write down solutions for $x(t)$, $y(t)$, $v_{x}(t)$, $v_{y}(t)$ in each of the three regions: I ($t\leq 0$), II($0\leq t\leq T$ and III($t\geq T$). The solutions for the coordinates and velocities have to be matched at the two boundaries at $t=0$ and $t=T$. This problem appears to have been discussed briefly in a very old paper gould , though the analysis there addresses a different motivation and purpose. We also know that the coupled equations can be rewritten using a complex velocity ${\bf\tilde{w}}=v_{x}+iv_{y}$ and solved directly using ${\bf\tilde{w}}={\bf\tilde{p}}\,e^{-i\omega_{B}t}$. The boundary conditions mentioned earlier will have to be imposed after writing down the solutions in each region.
Going through this rather straightforward exercise, we obtain all solutions in the three regions. These are given below. Here, $\omega_{B}=\frac{qB}{m}$ is the so–called cyclotron frequency, and the superscripts ‘I’, ‘II’ and ‘III’ refer to the region under consideration, as mentioned above. Arbitrary integration constants appear which will be fixed through chosen conditions on the coordinates and velocities. Note that there are four constants which can be fixed through conditions on position and velocity. We will, for now, keep them unspecified in the expressions below. While discussing examples and showing plots later, we will choose representative values.
Region I ($t\leq 0$):
$x^{(I)}=A_{0}\sin\,\omega_{B}t-D_{0}\cos\,\omega_{B}t+C_{0}$ (7)
$y^{(I)}=A_{0}\cos\,\omega_{B}t+D_{0}\sin\,\omega_{B}t+C^{\prime}_{0}$ (8)
$v_{x}^{(I)}=\omega_{B}A_{0}\cos\,\omega_{B}t+\omega_{B}D_{0}\sin\,\omega_{B}t$ (9)
$v_{y}^{(I)}=-\omega_{B}A_{0}\sin\,\omega_{B}t+\omega_{B}D_{0}\cos\,\omega_{B}t$ (10)
Region II ($0\leq t\leq T$):
$x^{(II)}=\left(\frac{E}{B}\sin\,\alpha\right)t+\left(A_{0}-\frac{E}{\omega_{B}B}\sin\,\alpha\right)\sin\,\omega_{B}t$
$-\left(D_{0}+\frac{E}{\omega_{B}B}\cos\,\alpha\right)\cos\,\omega_{B}t+\left(C_{0}+\frac{E}{\omega_{B}B}\cos\,\alpha\right)$ (11)
$y^{(II)}=-\left(\frac{E}{B}\cos\,\alpha\right)t+\left(A_{0}-\frac{E}{\omega_{B}B}\sin\,\alpha\right)\cos\,\omega_{B}t$
$+\left(D_{0}+\frac{E}{\omega_{B}B}\cos\,\alpha\right)\sin\,\omega_{B}t+\left(C^{\prime}_{0}+\frac{E}{\omega_{B}B}\sin\,\alpha\right)$ (12)
$v_{x}^{(II)}=\left(\frac{E}{B}\sin\,\alpha\right)+\omega_{B}\left(A_{0}-\frac{E}{\omega_{B}B}\sin\,\alpha\right)\cos\,\omega_{B}t$
$+\omega_{B}\left(D_{0}+\frac{E}{\omega_{B}B}\cos\,\alpha\right)\sin\,\omega_{B}t$ (13)
$v_{y}^{(II)}=-\left(\frac{E}{B}\cos\,\alpha\right)-\omega_{B}\left(A_{0}-\frac{E}{\omega_{B}B}\sin\,\alpha\right)\sin\,\omega_{B}t$
$+\omega_{B}\left(D_{0}+\frac{E}{\omega_{B}B}\cos\,\alpha\right)\cos\,\omega_{B}t$ (14)
Region III ($t\geq T$):
$x^{(III)}=\left[A_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\sin\,\omega_{B}t$
$-\left[D_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\cos\,\omega_{B}t$
$+\left(C_{0}+\frac{E}{B}\left(\sin\,\alpha\right)T\right)$ (15)
$y^{(III)}=\left[A_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\cos\,\omega_{B}t$
$+\left[D_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\sin\,\omega_{B}t$
$+\left(C^{\prime}_{0}-\frac{E}{B}\left(\cos\,\alpha\right)T\right)$ (16)
$v_{x}^{(III)}=\omega_{B}\left[A_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\cos\,\omega_{B}t$
$+\omega_{B}\left[D_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\sin\,\omega_{B}t$ (17)
$v_{y}^{(III)}=-\omega_{B}\left[A_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\sin\,\omega_{B}t$
$+\omega_{B}\left[D_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]\cos\,\omega_{B}t$ (18)
One may verify that the above coordinate solutions and velocity components match at the boundaries at $t=0$ and $t=T$. The time derivative of the velocity components are however discontinuous at $t=0$ and $t=T$ because of the discontinuous nature of the electric pulse, a fact evident from Eqns. (3), (4) and (5).
Figure 2: The circular paths before and after the arrival of the pulse. The blue (red) curve traces the motion of the charged particle before (after) the electric pulse is injected. Parameter values chosen (in respective units) are: $A_{0}=3$, $D_{0}=4$, $C_{0}=1$, $C^{\prime}_{0}=2$, $\omega_{B}=1$, $\alpha=2\pi$, $T=2$, $\frac{E}{B}=0.8$. The centre of the blue curve is at $(1,2)$ and its radius is 5 units. For the red curve, the centre is at $(1,0.4)$ and the radius is $6.34$ units.
Figure 3: The circular paths before and after the arrival of the pulse. The blue (red) curve traces the motion of the charged particle before (after) the electric pulse is injected. Parameter values chosen (in respective units) are: $A_{0}=-0.6$, $D_{0}=-1.2$, $C_{0}=1$, $C^{\prime}_{0}=2$, $\omega_{B}=1$, $\alpha=2\pi$, $T=2$, $\frac{E}{B}=0.8$. The centre of the blue curve is at $(1,2)$ and its radius is 1.342 units. For the red curve, the centre is at $(1,0.4)$ and the radius is $0.144$ units. In this case, the radius is smaller.
In Region I, one can clearly see that we have a circular trajectory given by the equation:
$\left(x^{(I)}-C_{0}\right)^{2}+\left(y^{(I)}-C^{\prime 0}\right)^{2}=A_{0}^{2}+D_{0}^{2}$ (19)
The centre of the circle is at $(C_{0},C^{\prime}_{0})$ and its radius is $\sqrt{A_{0}^{2}+D_{0}^{2}}$.
The effect of the pulse is manifest in the nature of the trajectories in regions II and III. In III, we find that the trajectory is still a circle but with a shifted centre and a different radius. The equation of this new circle at times $t\geq T$ is given as:
$\left(x^{(III)}-\left(C_{0}+\frac{E}{B}\left(\sin\,\alpha\right)T\right)\right)^{2}+\left(y^{(III)}-\left(C^{\prime}_{0}-\frac{E}{B}\left(\cos\,\alpha\right)T\right)\right)^{2}=$
$\left[A_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}+\left[D_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}$ (20)
Figure 4: Evolution of the charge in time. The red curve (pre-pulse circle) is from $t=-7$ to $t=0$ (clockwise). The blue curve is from $t=0$ to $t=5$ and the green curve (post-pulse circle) is from $t=5$ to $t=15$ (clockwise). The points $t=0$ and $t=5$ are the locations where the red, blue and the blue, green curves meet tangentially, as shown. Chosen parameters are: $\omega_{B}=1$, $T=5$, $\alpha=2\pi$, $A_{0}=1$, $D_{0}=2$, $C_{0}=3$ and $C^{\prime}_{0}=4$, $\frac{E}{\omega_{B}B}=3$.
Therefore, the centre of the circle now appears to be at $\left(C_{0}+\frac{E}{B}\left(\sin\,\alpha\right)T,C^{\prime}_{0}-\frac{E}{B}\left(\cos\,\alpha\right)T\right)$. Its radius is now changed (increased/decreased) to a a value equal to the square root of the expression given in the R. H. S. of Eqn. (20). It is easy to check and worth noting that there is no choice of values of any of the constants for which both the location of the centre and the radius remain unchanged.
The two circular paths before (blue curve) and after (red curve) the pulse are illustrated (using specific values of the parameters and constants) in Figures 2 and 3. The shift of the centre and the increase/decrease in the radius are clearly visible in both figures.
An observer watching the charge from a time much before $t=0$ will find it moving along a circular path of a certain radius. As one approaches the time $t=0$, the charge jumps to the trajectory (not closed) as specified in Eqns. (11) and (12). When $t=T$ is reached, the trajectory joins up with a circle with a different centre and a different radius. Subsequently, the charge is found to move on this new circular path forever. Figure 4 shows the evolution – the initial circle in red, the curved path in-between in blue and the final circular path in green. The shift of the centre of the circle and the change in the radius are once again manifest here, clearly. One may extend the time to values before $t=-7$ units and beyond $t=15$ units to any extent – in such a case we will just observe winding around the red and green circles along which the charge moves in the remote past and distant future, respectively.
Let us now note a few features and consequences of the main results stated above.
$\bullet$ With properly tuned initial conditions one may end up with $x,y$ as constant. This will happen when the values of $A_{0}$ and $D_{0}$ are chosen to be numerically given by:
$A_{0}=-\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)$ (21)
$D_{0}=-\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)$ (22)
As an example, for the parameter values $C_{0}=1$, $C^{\prime}_{0}=2$, $\omega_{B}=1$, $\alpha=2\pi$, $T=2$, $\frac{E}{B}=0.8$ (used to generate the curves in Figure 2), the choice of $A_{0}=-0.727$ and $D_{0}=-1.133$ yields a particle stuck and static at $(1,0.4)$ after the departure of the pulse. Note that the centre of the circular trajectory was initially at $(1,2)$. Figure 5 shows this result.
$\bullet$ For special values of $\omega_{B}T$ we have the following scenarios. When $\omega_{B}T=2n\pi$ ($n=0,1,2...$), the radius of the circle does not change, though its centre shifts. If, $\omega_{B}T=(2n+1)\pi$ ($n=0,1,2,...$), the radius changes according to the chosen value of $\alpha$. For $\omega_{B}T=-2\alpha$, the radius always decreases.
Figure 5: The circular path (green) before the arrival of the pulse. The blue curve traces the motion of the charged particle during the period when the electric pulse is present. The red dot at the end of the blue curve shows the static particle after the pulse has departed. Parameter values chosen (in respective units) are: $A_{0}=-0.727$, $D_{0}=-1.133$, $C_{0}=1$, $C^{\prime}_{0}=2$, $\omega_{B}=1$, $\alpha=2\pi$, $T=2$, $\frac{E}{B}=0.8$. The centre of the green curve is at $(1,2)$ and its radius is 1.342 units. The location (red dot) of the static particle at late times is at $(1,0.4)$.
$\bullet$ The change in the velocity components can be found by simply evaluating $\Delta v_{\alpha}=v_{\alpha}(t_{f})-v_{\alpha}(t_{i})$, where $t_{i}<0$ and $t_{f}>T$ and $\alpha\equiv x,y$. For no pair of values of $t_{i}$ and $t_{f}$ is the difference $\Delta v_{\alpha}=0$. This ensures the presence of a memory-like effect.
$\bullet$ One can also evaluate $v^{2}$ at any time $t_{i}<0$ and $t_{f}>0$. The difference $v_{f}^{2}-v_{i}^{2}$ turns out to be
$v_{f}^{2}-v_{i}^{2}=\omega_{B}^{2}\left\{\left[A_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}\right.$
$\left.+\left[D_{0}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}-A_{0}^{2}-D_{0}^{2}\right\}$ (23)
This, once again shows up as a memory-like effect (similar, but not exactly the same as the so-called ‘velocity kick’ biegarf1 ) induced by the electric field pulse.
$\bullet$ Let us now consider a pair of identical charges starting out with different initial conditions. We make the following choices for the constants:
Charge 1: $A_{01}$, $D_{01}$, $C_{01}$, $C^{\prime}_{01}$   ;    Charge 2: $A_{02}$, $D_{02}$, $C_{02}$, $C^{\prime}_{02}$.
The radii of the circular orbits before the charges encounter the pulse are $R_{i1}=\sqrt{A_{01}^{2}+D_{01}^{2}}$ and $R_{i2}=\sqrt{A_{02}^{2}+D_{02}^{2}}$. Hence the difference in the magnitudes of the radii and velocities are given as:
$\Delta R_{i}=|R_{i2}-R_{i1}|$    ,   $\Delta v_{i}=|v_{i2}-v_{i1}|=\omega_{B}\Delta R_{i}$.
Figure 6: The left figure shows the initial circular trajectories (before the arrival of the pulse) of radii $R_{i1}$ (blue) and $R_{i2}$ (yellow). The right figure shows the final circular trajectories (after the pulse has departed) of radii $R_{f1}$ (blue) and $R_{f2}$ (yellow). The chosen values are $A_{01}=D_{01}=\frac{1}{\sqrt{2}}$, $A_{02}=D_{02}=\sqrt{2}$, $C_{01}=C_{02}=1$, $C^{\prime}_{01}=C^{\prime}_{02}=2$, $\omega_{B}=1$, $T=\pi$, $\frac{E}{B}=0.5$, $\alpha=\frac{\pi}{4}$. Note that $\Delta R_{i}\neq\Delta R_{f}$ and hence $\Delta v_{i}\neq\Delta v_{f}$. In the graphs $\Delta R_{i}=1$, $\Delta R_{f}=0.8219$. The concentric centre of the two particles is initially at $(1,2)$. After the pulse departs it shifts to $(2.11,0.89)$.
After the pulse has interacted and departed, we have different radii and velocities. In particular, we have,
$R_{f1}^{2}=\left[A_{01}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}+$
$\left[D_{01}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}.$ (24)
$R_{f2}^{2}=\left[A_{02}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\cos\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}+$
$\left[D_{02}+\frac{2E}{\omega_{B}B}\sin\,\frac{\omega_{B}T}{2}\sin\left(\frac{\omega_{B}T}{2}+\alpha\right)\right]^{2}.$ (25)
Therefore,
$\Delta R_{f}=|R_{f2}-R_{f1}|$    ,   $\Delta v_{f}=|v_{f2}-v_{f1}|=\omega_{B}\Delta R_{f}$.
From the expressions above it is clear that $\Delta R_{i}\neq\Delta R_{f}$ and hence $\Delta v_{i}\neq\Delta v_{f}$. In addition the centre also shifts. This is shown in Figure 6. It also demonstrates the ‘permanent change’ character of the memory effect.
$\bullet$ The velocity field before and after the pulse have ${\bf\nabla\cdot v=0}$ and ×${\bf\nabla\times v\neq 0}$. Hence, one would say that the gradient of the velocity field has zero expansion and shear and nonzero rotation both before and after the electric pulse acts. The rotation is just proportional to $\omega_{B}$. One cannot therefore derive any link to a memory effect from the kinematics of a family of trajectories skrsadg . In other words there is no ${\cal B}$ memory olough , ic1 to be seen here.
$\bullet$ It may be possible to see the above-analysed memory effect in a cyclotron. Recall that one uses a radio-frequency alternating electric field (equal to $\omega_{B}$ and spread over a small region along the diameter of the dees) between the dees of a cyclotron, in order to drive the electron along a spiral path in the dees. In this process, the electron is also accelerated, which is indeed the purpose in a cyclotron. One will have to modify the set-up a bit so that one has a homogeneous, short-duration electric field apart from the magnets which give the uniform magnetic field. It may then be possible to see the change in the circular trajectory, caused by the pulse – i.e. see the change in radius, the shift of the centre and the change in the velocity. The practical implementation of this idea in a possible experiment will surely require more detailed thought which is left for the interested experimentalist to worry about.
$\bullet$ In order to get a better picture, let us now find some numbers. We assume $m=3.3\times{10}^{-27}$ kg (deuteron) , $q=1.6\times{10}^{-19}$ C and $B=1.5$ Tesla. $\alpha$ is chosen to be equal to $2\pi$. These values imply $\omega_{B}=72.7\times{10}^{6}$ rad/s which is in the radio frequency range (around 12 MHz). We also assume $T=10$ nanoseconds and $E=50$ MegaVolts/m. The constants are chosen as $A_{0}=0.30$ m, $D_{0}=0.40$ m, $C_{0}=0.60$ m and $C^{\prime}_{0}=-0.10$ m. The chosen values yield the following:
Initial trajectory: Centre at $(0.60,-0.10)$, Radius $R_{i}=0.50$ m and $v_{i}=0.12$ c.
Final trajectory: Centre at $(0.60,-0.43)$, Radius $R_{f}=0.79$ m and $v_{f}=0.19$ c.
From these numbers one may get a quantitative idea of the amount by which the centre may shift and how much the radius and velocity can change due to a short duration electric pulse. Note that all the velocities are non-relativistic in value.
$\bullet$ It is possible to write down the current ${\bf J}$ which produces the electric pulse using the Ampere-Maxwell law. The electric pulse can be written as
${\bf E}=E\left[\Theta(t)-\Theta(t-T)\right]\left(\cos\,\alpha\,{\bf\hat{i}}+\sin\,\alpha\,{\bf\hat{j}}\right).$ (26)
This gives a current ${\bf J}$ (in vacuum) equal to
${\bf J}=\epsilon_{0}E\left[\delta(t)-\delta(t-T)\right]\left(\cos\,\alpha\,{\bf\hat{i}}+\sin\,\alpha\,{\bf\hat{j}}\right).$ (27)
The current appears as a spike at $t=0$, creates the uniform electric field which is sustained till $t=T$. At $t=T$, there is a current spike along the opposite direction which reduces the net electric field to zero. This is how the short duration pulse seems to act.
$\bullet$ One may write down a formula analogous to Eqn. (1) for the system discussed here. Recall the complex velocity ${\bf\tilde{w}}=v_{x}+iv_{y}={\bf\tilde{p}}\,e^{-i\omega_{B}t}$. Solving the differential equation for ${\bf\tilde{w}}$, we can formally write
[…]
One can find the ‘velocity kick’ from this expression and it generalises Eqn. (1) when both electric and magnetic fields are present. In principle, it is possible to replace the electric field in terms of the current density vector (integral over it) and write down the ‘velocity kick’ solely in terms of the source properties, as has been the usual practice in memory calculations.
III Conclusions
In this brief article, we have shown how an effect somewhat similar to what is conventionally termed as electromagnetic memory can arise in a rather simple scenario. The age-old problem of a charged particle moving on a circular path due to a uniform magnetic field is modified by introducing a short-duration electric pulse. This causes a ‘permanent change’–the radius can increase/decrease, the centre of the circle always shifts away and the velocity as well as its magnitude can increase or decrease. We also showed that for a pair of particles moving in the same ${\bf B}$ and ${\bf E}$ fields but with different conditions at a given time, the relative separation, velocity do change. This is analogous to memory as found using a gravitational wave pulse where, a change in relative separation distinguishes between pulses ‘with memory’ or ‘without memory’. We have also mentioned (with some explicit values) how our results may be seen in a cyclotron with necessary practical adjustments.
As mentioned in the Introduction, it is true that the example discussed here does not concern the usual ‘wave memory’. It is, rather, a completely non-relativistic scenario where the effect of a short duration pulse on a trajectory is explored. One may argue that this can indeed also happen in simple mechanical systems – say, when a particle is acted upon by an impulsive force. In particle mechanics though, the particle will actually have to be ‘struck’ in order to give an impulse. This is not the case when a charged particle moves in magnetic and electric fields–the interaction of the charge with the magnetic and electric field is not a ‘contact’ interaction and may be visualised as an ‘influence’ over a short duration, which leads to a change which persists. Even though we do not have a ‘wave field’ here, the ‘time dependent pulse’ can bring about a change akin to memory. Such a pulse-induced memory is not uncommon in condensed matter (magnetic) systems as is well-known and well-studied in the literature (see eg. jacs , acs among others).
An obvious extension of this work is to try out different pulse shapes –delta function, triangular, Gaussian, sech-squared are the simplest examples –the last two being smooth functions. It may be difficult to do the smooth function examples analytically, but all scenarios are surely tractable numerically. An important related question is about whether there is a pulse for which we can observe a shape change in the trajectory–i.e. a circle going over to an ellipse or something else. Such a model will be useful because it is, in some sense, perhaps closer to the idea of memory as proposed and pursued in the context of gravitational wave physics ayansayan .
ACKNOWLEDGEMENTS
The author thanks Indranil Chakraborty for his comments and suggestions on the manuscript.